OpenStudy (anonymous):

need help solving \[\sqrt{3}(\sqrt{48}-8\sqrt{3})\]

OpenStudy (gorv):

\[\sqrt{3}*(\sqrt{16*3}-8*\sqrt{3})=\sqrt{3}*(4\sqrt{3}-8\sqrt{3})=\sqrt{3}(-4\sqrt{3})=-4*\sqrt{3}*\sqrt{3}=-4*3=-12\]

OpenStudy (gorv):

\[-4*\sqrt{3}*\sqrt{3}=-4*3=-12\]

OpenStudy (anonymous):

thank you

OpenStudy (gorv):

welcome :)